### v1

# An extended NJL model for baryonic matter and quark matter

Abstract
By considering baryons as clusters of three quarks, we extend the Nambu-Jona-Lasinio (NJL) model to describe baryonic matter, quark matter, and their transitions in a unified manner, where the Dirac sea and spontaneous chiral symmetry breaking are considered. In particular, a density-dependent structural function $\alpha_{S}$ is introduced to modulate the four(six)-point interaction strengths to reproduce the baryon masses in vacuum and medium. Vector interactions are considered with the exchange of $\omega$ and $\rho$ mesons, where the density-dependent coupling constants are fixed by reproducing nuclear matter properties as well as the $\Lambda$-hyperon potential depth in nuclear medium. As density increases, quarks will emerge as quasi-free particles and coexist with baryons. This phase is interpreted as quarkyonic matter, where quarks are restricted to the lowest energy states in the presence of baryons, i.e., quarks are still confined. Similar to the treatment of $\alpha$ clustering in nuclear medium, a Pauli blocking term is added to baryon masses so that baryons eventually become unbound in the presence of a quark Fermi Sea. Then at large enough densities baryons vanish and a Mott transition takes place, we consider such a transition as deconfinement phase transition. Depending on the strengths of Pauli blocking term and quark-vector meson couplings, both first-order and continues phase transitions are observed for quarkyonic, chiral, and deconfinement phase transitions. The corresponding compact star structures are then investigated and confronted with various astrophysical constraints.
I Introduction
Cold strongly interacting matter are expected to undergo a deconfinement phase transition at large enough densities, converting hadronic matter into quark matter. Due to the infamous sign problem in lattice QCD and the inability to carry out reliable pertubative QCD calculations, so far it is still unclear at which density does deconfinement phase transition take place or whether it is of first order or a smooth crossover Fukushima et al. (2005); Voskresensky (2023). Such a situation may be improved in the era of multi-messenger astronomy, where recent Bayesian analysis Xie and Li (2021a); Annala et al. (2023); Han et al. (2023); Pang et al. (2024) and binary neutron star merger simulations Bauswein et al. (2019); Huang et al. (2022); Fujimoto et al. (2023) have shed light on identifying a possible deconfinement phase transition in massive compact stars.
For a first-order deconfinement phase transition Dexheimer and Schramm (2010), the properties of hadronic matter and quark matter are distinctively different. A quark-hadron mixed phase is thus expected in hybrid stars, which may exhibit various geometrical structures based on the surface tension of quark-hadron interface and consequently impact hybrid star structures Heiselberg et al. (1993); Glendenning (2001); Voskresensky et al. (2002); Tatsumi et al. (2003); Voskresensky et al. (2003); Endo et al. (2005); Maruyama et al. (2007); Peng et al. (2008); Yasutake et al. (2014); Maslov et al. (2019); Xia et al. (2019, 2020). Alternatively, the hadron-quark crossover at finite densities was proposed, which predicts a stiff equation of state (EOS) to support massive hybrid stars Baym (1979); Çelik et al. (1980); Schäfer and Wilczek (1999); Fukushima (2004); Hatsuda et al. (2006); Maeda et al. (2009); Masuda et al. (2013a, b); Zhao et al. (2015); Kojo et al. (2015); Masuda et al. (2016); Whittenbury et al. (2016); Bai et al. (2018); Baym et al. (2019). As demonstrated by Fukushima and Kojo Fukushima and Kojo (2016), the crossover from hadronic matter to quark matter can be bridged by quarkyonic matter, which is comprised of “a quark Fermi Sea” and “a baryonic Fermi surface” McLerran (2009); McLerran and Pisarski (2007); McLerran et al. (2009).
At densities below and around the nuclear saturation density $n_{0}\approx 0.16$ fm-3, it was shown that treating nucleons as the basic building blocks gives satisfactory description for finite nuclei and nuclear matter, e.g., relativistic mean field (RMF) models Meng (2016), nonrelativistic density functional methods Dutra et al. (2012), and microscopic many-body theories Hebeler (2021). However, in nuclear medium the effects of quark degrees of freedom will emerge, such as the EMC effect with the nuclear structure functions differ from those of free nucleons European Muon Collaboration (1983). By treating nucleons as three quarks confined in bags with the interaction mediated by mesons, the Quark-meson-Coupling (QMC) model was proposed, where the EMC effect in the valence region could be understood Guichon et al. (2018).
When density becomes larger, it is nevertheless insufficient to treat baryons as independent particles, where the number of exchanged quarks between baryons increase and baryons may overlap with each other Fukushima and Kojo (2016); McLerran (2009); Park et al. (2021); Park and Lee (2022). It was shown that quark wave functions simultaneously exist in different baryons to support the attractive interactions in the intermediate range Chen et al. (2007), while the short range repulsion is provided by the Pauli principle Oka and Yazaki (1980). In such cases, to unveil the microscopic dynamics of baryonic matter, quark matter, and their transitions, a unified description based on quark degrees of freedom is essential, where extensive efforts were made in the past decades. For example, Horowitz and Piekarewicz investigated the hadron-quark transition based on a nonrelativistic one dimensional string-flip model Horowitz and Piekarewicz (1991). A microscopic approach of color molecular dynamics was proposed by Maruyama and Hatsuda Maruyama and Hatsuda (2000), where quarks are clusterized into baryons at low density and a deconfined quark matter is formed at high density via crossover Yasutake and Maruyama (2024). By treating nucleon as a quark-diquark state using the Nambu-Jona-Lasinio (NJL) model, it was shown that the effects of confinement leads to a scalar polarizability of the nucleon and a less attractive effective interaction between nucleons, helping to achieve saturation of the nuclear matter ground state Bentz and Thomas (2001). The hadronic SU(3) nonlinear $\sigma$ model was extended to quark degrees of freedom, suggesting that the deconfinement phase transition is of first order Dexheimer and Schramm (2010). A direct modeling of quarkyonic matter was recently proposed, where the pressure and sound velocity increase rapidly with density McLerran and Reddy (2019); Margueron et al. (2021). By synthesizing the Walecka model and quark-meson model, a complete field model was developed Cao and Liao (2020); Cao (2022), which well describes the chiral symmetry restoration in quarkyonic matter.
In our previous study, by combining RMF models and equivparticle models with density-dependent quark masses, we have investigated nuclear matter, quarkyonic matter, and quark matter in a unified manner Xia et al. ; Xia et al. (2023). Nevertheless, the dynamic chiral symmetry restoration was not addressed, and the contribution of Dirac sea was neglected. To better describe strongly interacting matter at vast density ranges, it is favorable to start from quark degrees of freedom directly and consider baryons as clusters made of three valence quarks. As density increases, similar to the melting of light clusters in nuclear medium Typel et al. (2010), the deconfinement phase transition can be viewed as a Mott transition of quark clusters Bastian et al. (2018). In this work, we thus extend the NJL model to describe baryonic matter, quark matter, and their transitions in a unified manner, where the Dirac sea and spontaneous chiral symmetry breaking are accounted for. To reproduce the baryon masses in vacuum and medium, the four(six)-point interaction strengths in NJL model are modulate by a density-dependent structural function. The repulsive interaction is treated with the exchange of $\omega$ and $\rho$ mesons, where density-dependent couplings are fixed by reproducing nuclear matter properties as well as the $\Lambda$-hyperon potential depth in nuclear medium.
The paper is organized as follows. In Section II, we present the theoretical framework for the extended NJL model with the model parameters fixed based on various constraints. Several types of phase transitions and the properties of dense stellar matter are then examined, where the corresponding compact star structures are confronted with various astrophysical constraints in Section III. We draw our conclusion in Section IV.
II Theoretical framework
II.1 Lagrangian density
In the mean-field approximation, the Lagrangian density of an extended SU(3) NJL model is given by
$\mathcal{L}$ $=$ $\sum_{i}\bar{\psi}_{i}\left(i\gamma_{\mu}D^{\mu}_{i}-M_{i}\right)\psi_{i}-\frac{1}{4}\omega_{\mu\nu}\omega^{\mu\nu}+\frac{1}{2}m_{\omega}^{2}\omega^{2}$ (1)
$\mbox{}+\frac{1}{2}\sum_{i=u,d,s}\left[\partial_{\mu}\sigma_{i}\partial^{\mu}\sigma_{i}-m_{\sigma}^{2}\sigma_{i}^{2}\right]-\frac{1}{4}\vec{\rho}_{\mu\nu}\cdot\vec{\rho}^{\mu\nu}$
$\mbox{}+\frac{1}{2}m_{\rho}^{2}\rho^{2}+4K\bar{n}_{u}^{s}\bar{n}_{d}^{s}\bar{n}_{s}^{s}.$
Here $\psi_{i}$ represents the Dirac spinor for different fermions $i$, where in this work we have included baryons ($p$, $n$, $\Lambda$), quarks ($u$, $d$, $s$) and leptons ($\mu$, $e$). By replacing the quark condensations with $\sigma_{u,d,s}$ fields, the four-point interaction becomes nonlocal, while the six-point (‘t Hooft) interaction remains local. In principle, the diquark coupling terms can also be included to treat pairings among baryons and quarks, which should be explored in our future study. Vector meson fields are also introduced to account for the repulsive interactions, where the field tensors are
$\omega_{\mu\nu}=\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu},\ \ \vec{\rho}_{\mu\nu}=\partial_{\mu}\vec{\rho}_{\nu}-\partial_{\nu}\vec{\rho}_{\mu}.$ (2)
The covariant derivative in Eq. (1) takes the form
$iD^{\mu}_{i}=i\partial^{\mu}-f_{i}g_{\omega}\sum_{q=u,d,s}N^{q}_{i}\omega^{\mu}-f_{i}g_{\rho}\vec{\tau}_{i}\cdot\vec{\rho}^{\mu},$ (3)
where $N^{q}_{i}$ is the number of valence quarks $q$ in particle $i$ and $\vec{\tau}_{i}$ the isospin. The factor $f_{i}$ modulates the coupling strengths between particle $i$ and vector mesons, while $g_{\omega}$ and $g_{\rho}$ are fixed according to nuclear matter properties with $f_{p}=f_{n}=1$. In this work, we consider baryons as clusters of quarks with their effective masses given by
$M_{i}=\sum_{q=u,d,s}N^{q}_{i}\left[m_{q0}+\alpha_{S}(M_{q}-m_{q0})\right]+Bn_{\mathrm{b}}^{Q}$ (4)
and quark masses by
$M_{q}=m_{q0}-g_{\sigma}\sigma_{q}+2K\frac{\bar{n}_{u}^{s}\bar{n}_{d}^{s}\bar{n}_{s}^{s}}{\bar{n}_{q}^{s}}$ (5)
with the effective quark scalar density
$\bar{n}_{q}^{s}=n_{q}^{s}+\alpha_{S}\sum_{i=p,n,\Lambda}N^{q}_{i}n_{i}^{s},$ (6)
where $n_{i}^{s}$ is fixed by Eq. (12). A density dependent structural function $\alpha_{S}$ is introduced for baryons, which mimics the dampened interaction strength as chiral condensates diminish within baryons Bentz and Thomas (2001); Reinhardt and Weigel (2012); Xia et al. (2014). The last term in Eq. (4) accounts for the effects of Pauli blocking and interactions between quarks and baryons with $n_{\mathrm{b}}^{Q}=(n_{u}+n_{d}+n_{s})/3$ being the baryon number density of quarks Xia et al. ; Xia et al. (2023), which resembles the treatments of $\alpha$ clustering inside nuclear matter Röpke et al. (2014); Xu et al. (2016). The total baryon number density is then
$n_{\mathrm{b}}=n_{p}+n_{n}+n_{\Lambda}+n_{\mathrm{b}}^{Q},$ (7)
where the number density of particle $i$ is determined by Eq. (11). The masses of leptons remain constant with $M_{e}=0.511~{}$MeV and $M_{\mu}=105.66~{}$MeV Particle Data Group (2014).
Due to time-reversal symmetry and charge conservation, the boson fields take mean values with only the time component and the $3$rd component in the isospin space. We then define $\omega\equiv\omega_{0}$, $\rho\equiv\rho_{0,3}$, and $\tau_{i}\equiv\tau_{i,3}$ for simplicity. Based on the Lagrangian density in Eq. (1), the meson fields of uniform dense matter are determined by
$m_{\sigma}^{2}\sigma_{q}$ $=$ $g_{\sigma}\bar{n}_{q}^{s}$ (8)
$m_{\omega}^{2}\omega$ $=$ $g_{\omega}\sum_{i}f_{i}\sum_{q=u,d,s}N^{q}_{i}n_{i},$ (9)
$m_{\rho}^{2}\rho$ $=$ $g_{\rho}\sum_{i}f_{i}\tau_{i}n_{i}.$ (10)
At vanishing temperatures, the number density and scalar density of fermion $i$ in uniform dense matter are
$n_{i}$ $=$ $\langle\bar{\psi}_{i}\gamma^{0}\psi_{i}\rangle=\frac{g_{i}\nu_{i}^{3}}{6\pi^{2}},$ (11)
$n_{i}^{s}$ $=$ $\langle\bar{\psi}_{i}\psi_{i}\rangle=\frac{g_{i}M_{i}^{3}}{4\pi^{2}}\left[x_{i}\sqrt{x_{i}^{2}+1}-\mathrm{arcsh}(x_{i})\right.$ (12)
$\left.-y_{i}\sqrt{y_{i}^{2}+1}+\mathrm{arcsh}(y_{i})\right].$
Here we take $x_{i}\equiv\nu_{i}/M_{i}$ with $\nu_{i}$ being the Fermi momentum, $y_{i}\equiv\Lambda/M_{i}$ with $\Lambda$ being the 3-momentum cut-off to regularize the vacuum part of quarks ($\Lambda=0$ for baryons and leptons), and the degeneracy factors $g_{n,p,\Lambda}=g_{e,\mu}=2$ and $g_{u,d,s}=6$. Note that the baryonic Dirac sea does not exist since quarks no longer form clusters in the Dirac sea due to Pauli blocking. Meanwhile, it is worth mentioning that in the quarkyonic phase with baryons and quarks coexisting in a same volume, only baryons can be exited to higher energy states while quarks are restricted to the lowest energy states as they are still confined.
The energy density of uniform dense matter is fixed by
$E$ $=$ $\sum_{i}\frac{g_{i}{M_{i}}^{4}}{16\pi^{2}}\left[x_{i}(2x_{i}^{2}+1)\sqrt{x_{i}^{2}+1}-\mathrm{arcsh}(x_{i})\right]$ (13)
$-\sum_{i=u,d,s}\frac{g_{i}{M_{i}}^{4}}{16\pi^{2}}\left[y_{i}(2y_{i}^{2}+1)\sqrt{y_{i}^{2}+1}-\mathrm{arcsh}(y_{i})\right]$
$+\frac{1}{2}\left[m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}+\sum_{i=u,d,s}m_{\sigma}^{2}\sigma_{i}^{2}\right]-4K\bar{n}_{u}^{s}\bar{n}_{d}^{s}\bar{n}_{s}^{s}$
$-{E}_{0}.$
Here a constant $E_{0}$ is introduced to ensure $E=0$ in the vacuum. The chemical potentials of baryon $b$, quark $q$, and lepton $l$ are fixed by
$\mu_{b}$ $=$ $\sqrt{\nu_{b}^{2}+M_{b}^{2}}+f_{b}(3g_{\omega}\omega+g_{\rho}\tau_{b}\rho)+\Sigma_{b}^{\mathrm{R}},$ (14)
$\mu_{q}$ $=$ $\sqrt{\nu_{q}^{2}+M_{q}^{2}}+g_{\omega}\omega+g_{\rho}\tau_{q}\rho+\Sigma_{q}^{\mathrm{R}},$ (15)
$\mu_{l}$ $=$ $\sqrt{\nu_{l}^{2}+M_{l}^{2}},$ (16)
with the “rearrangement” terms given by
$\Sigma_{b}^{\mathrm{R}}$ $=$ $\sum_{i}f_{i}\left(\omega n_{i}\sum_{q=u,d,s}N^{q}_{i}\frac{\mbox{d}g_{\omega}}{\mbox{d}n_{\mathrm{b}}}+\rho\tau_{i}n_{i}\frac{\mbox{d}g_{\rho}}{\mbox{d}n_{\mathrm{b}}}\right)$ (17)
$+\sum_{i=n,p,\Lambda}\left[\frac{\mbox{d}\alpha_{S}}{\mbox{d}n_{\mathrm{b}}}\sum_{q=u,d,s}N^{q}_{i}(M_{q}-m_{q0})\right]n_{i}^{s},$
$\Sigma_{q}^{\mathrm{R}}$ $=$ $\frac{1}{3}B\sum_{i=n,p,\Lambda}n_{i}^{s}+\frac{1}{3}\Sigma_{b}^{\mathrm{R}}.$ (18)
Then the pressure $P$ is obtained with
$P=\sum_{i}\mu_{i}n_{i}-E.$ (19)
II.2 Model parameters
In this work we adopt the RKH parameter set of NJL model, i.e., $\Lambda=602.3$ MeV, $m_{u0}=m_{d0}=5.5$ MeV, $m_{s0}=140.7$ MeV, $G_{S}=g_{\sigma}^{2}/4m_{\sigma}^{2}=1.835/\Lambda^{2}$, and $K=12.36/\Lambda^{5}$ Rehberg et al. (1996). This parameter set well reproduces the meson properties, i.e., $m_{\pi}=135.0$ MeV, $m_{K}=497.7$ MeV, $m_{\eta^{\prime}}=957.8$ MeV, $m_{\eta}=514.$ MeV, and $f_{\pi}=92.4$ MeV Rehberg et al. (1996). The corresponding vacuum quark condensates are $\bar{n}_{u}^{s}=\bar{n}_{d}^{s}=-1.843\ \mathrm{fm}^{-3}$ and $\bar{n}_{s}^{s}=-2.227\ \mathrm{fm}^{-3}$, indicating the quark masses $M_{u}=M_{d}=367.6\ \mathrm{MeV}$ and $M_{s}=549.5\ \mathrm{MeV}$.
To reproduce the masses of nucleons and $\Lambda$-hyperons in vacuum, we take $\alpha_{S}=0.849$ at $n_{\mathrm{b}}=0$, which gives $M_{p}=M_{n}=938.9$ MeV and $M_{\Lambda}=1113.7$ MeV by applying Eq. (4). Meanwhile, as was done in traditional RMF models Meng (2016), in addition to the repulsive interaction exerted by vector mesons, we take $\alpha_{S}=0.57$ at $n_{\mathrm{b}}=n_{S}=0.16\ \mathrm{fm}^{-3}$ so that a strong attractive interaction ($M_{p}=M_{n}=519.2$ MeV) is present to accommodate the spin-orbit splittings in finite nuclei, which is essential to reproduce the nuclear magic numbers. We then write out the following formula
$\alpha_{S}=a_{S}\exp(-n_{\mathrm{b}}/n_{S})+b_{S},$ (20)
which reproduce the values $\alpha_{S}(0)=0.849$ and $\alpha_{S}(n_{S})=0.57$ with $a_{S}$, $n_{S}$ and $b_{S}$ presented in Table 1. In Fig. 1 we present the masses of nucleons, $\Lambda$-hyperons, and quarks in symmetric nuclear matter (SNM), which are obtained with Eqs. (4) and (5). Evidently, the masses of nucleons and $\Lambda$-hyperons decrease quickly at $n_{\mathrm{b}}\lesssim n_{S}$ due to the structural function (20). At larger densities, the masses of $u$-quarks, $d$-quarks, and nucleons further decrease and approach to their current masses with vanishing quark condensates, while the $s$-quark condensate $\bar{n}_{s}^{s}$ remains large at $n_{\mathrm{b}}\lesssim 1\ \mathrm{fm}^{-3}$ so that the masses of $s$-quarks and $\Lambda$-hyperons are still sizable in Fig. 1.
Figure 1: Masses of nucleons, $\Lambda$-hyperons, and quarks in symmetric nuclear matter (SNM) as functions of the total baryon number density $n_{\mathrm{b}}$, which are determined by adopting the density dependent structural function (20).
Table 1: The adopted parameter sets for the density dependent coupling constants in Eqs. (20-22).
$a_{S}=0.4413715$ $n_{S}=0.16\ \mathrm{fm}^{-3}$ $b_{S}=0.4076285$
$a_{V}=3.4618068$ $n_{V}=0.214\ \mathrm{fm}^{-3}$ $b_{V}=1.1067389$
$a_{TV}=0.5014459$ $n_{TV}=0.1\ \mathrm{fm}^{-3}$ $b_{TV}=0.0117601$
Figure 2: Energy per baryon $E/n_{\mathrm{b}}$ of nuclear matter as functions of the total baryon number density $n_{\mathrm{b}}$, which are obtained adopting the parameter sets indicated in Table. 1.
To fix the coupling constants of vector mesons, we adopt the well-constrained nuclear matter properties at $n_{0}=0.16\ \mathrm{fm}^{-3}$ and $n_{\mathrm{on}}=0.1\ \mathrm{fm}^{-3}$, i.e., $\varepsilon(n_{0})=-16$ MeV, $S(n_{0})=31.7\pm 3.2$ MeV Li and Han (2013); Oertel et al. (2017), $\varepsilon(n_{\mathrm{on}})=-14.1\pm 0.1$ MeV, and $S(n_{\mathrm{on}})=25.5\pm 1.0$ MeV Centelles et al. (2009); Brown (2013). In practice, with the nucleon masses being $m_{N}=M_{p}(0)=M_{n}(0)=938.9$ MeV, we reproduce the energy per baryon $E(n_{\mathrm{on}})/n_{\mathrm{on}}=m_{N}+\varepsilon(n_{\mathrm{on}})=924.8$ MeV and $E(n_{0})/n_{0}=m_{N}+\varepsilon(n_{0})=922.9$ MeV for SNM and $E(n_{\mathrm{on}})/n_{\mathrm{on}}=m_{N}+\varepsilon(n_{\mathrm{on}})+S(n_{\mathrm{on}})=950.3$ MeV and $E(n_{0})/n_{0}=m_{N}+\varepsilon(n_{0})+S(n_{0})=954.6$ MeV for pure neutron matter (PNM) adopting the following density dependent coupling constants, i.e.,
$g_{\omega}^{2}/m_{\omega}^{2}$ $=$ $4G_{S}[a_{V}\exp(-n_{\mathrm{b}}/n_{V})+b_{V}],$ (21)
$g_{\rho}^{2}/m_{\rho}^{2}$ $=$ $4G_{S}[a_{TV}\exp(-n_{\mathrm{b}}/n_{TV})+b_{TV}].$ (22)
The coefficients in Eqs. (21) and (22) used in this work are indicated in Table 1. Note that the meson masses $m_{\sigma}$, $m_{\omega}$, and $m_{\rho}$ are left undetermined in this work, which should be fixed in our future study according to the properties of finite nuclei.
Table 2: The saturation properties of nuclear matter predicted by the extended NJL model introduced in this work.
$n_{0}$ $\varepsilon$ $K$ $S$ $L$
fm-3 MeV MeV MeV MeV
0.158 $-16.0$ 234.5 31.5 42.4
In Fig. 2 we then present the energy per baryon of SNM and PNM, which are obtained with Eq. (13) adopting the parameter sets indicated in Table. 1. Evidently, the binding energies of both SNM and PNM reproduce the central values of the binding energies $\varepsilon(n_{0})=-16$ MeV, $\varepsilon(n_{\mathrm{on}})=-14.1\pm 0.1$ MeV and symmetry energies $S(n_{0})=31.7\pm 3.2$ MeV, $S(n_{\mathrm{on}})=25.5\pm 1.0$ MeV Li and Han (2013); Oertel et al. (2017); Centelles et al. (2009); Brown (2013). The corresponding saturation properties such as the incompressibility $K$ and slope $L$ of nuclear symmetry energy are indicated in Table 2, which are consistent with the state-of-art constraints $K=240\pm 20$ MeV Shlomo et al. (2006) and $L=58.7\pm 28.1$ MeV Li and Han (2013); Oertel et al. (2017). Note that constraints on higher order terms of nuclear saturation properties can be attained based astrophysical observations, heavy-ion collisions, measurements of the neutron skin thicknesses, and nuclear theories Zhang et al. (2020); Xie and Li (2021b); PREX Collaboration (2021); Essick et al. (2021); CREX Collaboration (2022).
Figure 3: Potential depth of $\Lambda$-hyperons in symmetric nuclear matter.
Beside nucleons, the contribution of $\Lambda$-hyperons is also considered in this work. Based on previous investigations of $\Lambda$-hypernuclei, it was shown that a $\Lambda$ potential well depth $U_{\Lambda}(n_{0})=-30$ MeV in SNM is required to accommodate the single $\Lambda$ binding energies Gal et al. (2016); Sun et al. (2016, 2018); Rong et al. (2021). In the framework of the extended NJL model, the potential depth $U_{\Lambda}(n_{0})=-30$ MeV indicates $f_{\Lambda}=1.0626$ for the vector meson couplings. The corresponding potential depth of $\Lambda$-hyperons is then fixed by $U_{\Lambda}(n_{\mathrm{b}})=\mu_{\Lambda}-M_{\Lambda}(0)$, where the corresponding values are presented in Fig. 3. The onset densities of $\Lambda$-hyperons in SNM and PNM are then $n_{\mathrm{onset}}=0.72$ and $0.44\ \mathrm{fm}^{-3}$, respectively.
III Results and discussions
Based on the formulae and model parameters presented in Sec. II, we then investigate the properties of dense stellar matter, where baryonic matter, quark matter, and quarkyonic matter can be treat in a unified manner. For compact star matter at a fixed total baryon number density $n_{\mathrm{b}}$, the densities of fermions are obtained with Eqs. (14-16) fulfilling the $\beta$-stability condition, i.e.,
$\mu_{i}=B_{i}\mu_{\mathrm{b}}-q_{i}\mu_{e},$ (23)
where $\mu_{\mathrm{b}}$ is the baryon chemical potential with $B_{i}$ ($B_{p}=B_{n}=B_{\Lambda}=1$, $B_{u}=B_{d}=B_{s}=1/3$, and $B_{e}=B_{\mu}=0$) being the baryon number of particle type $i$. Due to charge screening, the charge neutrality condition needs to be satisfied as well, i.e.,
$\sum_{i}q_{i}n_{i}\equiv 0,$ (24)
where $q_{n}=0$, $q_{p}=1$, $q_{u}=2/3$, $q_{d}=-1/3$, and $q_{e}=q_{\mu}=-1$ are the charge number of each particle type. The EOSs of compact star matter are obtained with the energy density $E$ fixed by Eq. (13) and pressure $P$ by Eq. (19).
Figure 4: Pressure $P$ as functions of baryon chemical potential $\mu_{\mathrm{b}}$ adopting various parameter sets with $f_{u}=f_{d}=f_{s}=0.5$, $B=0$ and 1 GeV/fm3. A first-order quarkyonic transition is observed if we take $B=0$, while first-order chiral and deconfinement transitions are identified if $B=1$ GeV/fm3.
We first examine the phase structure of dense stellar matter taking $f_{u}=f_{d}=f_{s}=0.5$, which approximately corresponds to the vector coupling $G_{V}\approx f_{q}b_{V}G_{S}=0.55G_{S}$ in traditional NJL models. Two scenarios for the Pauli blocking term in Eq. (4) are considered, i.e., $B=0$ and 1 GeV/fm3. For uniform compact star matter, the thermodynamical potential density is $\Omega=-P$. The stable phase at a fixed baryon chemical potential $\mu_{\mathrm{b}}$ is fixed by minimizing the thermodynamical potential density $\Omega$ or maximizing the pressure $P$.
In Fig. 4 we present the obtained pressure as functions of baryon chemical potential. Various first-order phase transitions are identified, where the solid stars indicate the coexistence points of those first-order phase transitions according to Maxwell construction of mixed phases. In principle, the Maxwell construction is valid only when the surface tension is large enough so that geometrical structures are unfavorable, which is often not the case Heiselberg et al. (1993); Voskresensky et al. (2002); Tatsumi et al. (2003); Voskresensky et al. (2003); Endo et al. (2005); Maruyama et al. (2007); Yasutake et al. (2014); Xia et al. (2019); Maslov et al. (2019); Xia et al. (2020). A more detailed investigation on the structures and properties of the mixed phases is thus necessary and should be carried out in our future study, while for now we restrict ourself to the Maxwell construction scenarios. In this work, we consider the following phase transitions:
  • Quarkyonic phase transition: emergence of quasi-free quarks in baryonic matter, where quarks are still confined and only baryons can be exited to higher energy states;
  • Chiral phase transition: quark condensate $\bar{n}_{q}^{s}$ ($q=u,d,s$) vanishes and quark mass $M_{q}$ is reduced to its current mass $m_{q0}$;
  • Deconfinement phase transition: at large enough densities baryons become unbound and start to dissolve, i.e., Mott transition.
All three types of phase transitions are identified in dense stellar matter, which become of first order if we adopt certain parameter sets. For example, as indicated in Fig. 4, the quarkyonic phase transition is of first order if we take $B=0$, which becomes of third order for $B=1$ GeV/fm3. Meanwhile, the chiral and deconfinement phase transitions become of first order only when we take large $B$, e.g., $B=1$ GeV/fm3.
Figure 5: Number densities of various fermions considered in this work. The yellow bands correspond to the density ranges of mixed phases indicated by the solid stars in Fig. 4, which are fixed by Maxwell construction.
The quarkyonic and deconfinement phase transitions are better illustrated in Fig. 5, where the number densities of various fermions are plotted. In particular, due to the requirement of charge neutrality, $d$-quarks emerge at $n_{\mathrm{b}}=n_{0}/2$ if we take $B=0$, which leads to a first-order phase transition with the reduction of energy per baryon. A mixed phase at $n_{\mathrm{b}}\approx 0.016$-0.24 fm-3 is then predicted by Maxwell construction, corresponding to the bulk separation of the quarkyonic and baryonic phases. If we adopt $B=1$ GeV/fm3, due to strong repulsive interactions between baryons and quarks, $d$-quarks emerge at much larger densities at $n_{\mathrm{b}}=0.17$ fm-3, while the quarkyonic phase transition becomes of third order. Similar scenarios are observed if we adopt larger $f_{q}$.
Meanwhile, it is found that the onset densities of $\Lambda$-hyperons are much larger with $n_{\mathrm{onset}}=0.51$ and 0.64 fm-3 when we take $B=0$ and 1 GeV/fm3, respectively. Due to the strong $\Lambda$-$\omega$ coupling, $\Lambda$-hyperons eventually vanish at larger densities. In such cases, $\Lambda$-hyperons have little impacts on the EOSs of compact star matter, so that the Hyperon Puzzle can be avoided Vidaña (2015).
As we further increase the density, baryons eventually become unbound and the corresponding number densities decrease quickly. A deconfinement phase transition then takes place as baryons vanish at $n_{\mathrm{b}}\approx 7$ (1.3) fm-3 when we take $B=0$ (1) GeV/fm3, which is of third (first) order. At largest densities, as indicated in Fig. 5, the quark number densities in strange quark matter become approximately the same with $n_{u}=n_{d}=n_{s}$, which favors the formation of color-flavor locked phase Rüster et al. (2005). This is nevertheless out of the scope of our current study and should be examined in future. Note that the density for deconfinement phase transition increases significantly as $f_{q}$ ($q=u,d,s$) approaches to 1.
Figure 6: Masses of nucleons, $\Lambda$-hyperons, and quarks in dense stellar matter as functions of the total baryon number density $n_{\mathrm{b}}$, corresponding to those in Figs. 4 and 5.
In between the quarkyonic and deconfinement phase transitions, a chiral phase transition takes place, where the quark condensates $\bar{n}_{u}^{s}$ and $\bar{n}_{d}^{s}$ vanish. It is found that the chiral phase transition is usually a smooth crossover when we take $B=0$, which becomes of first order if $B=1$ GeV/fm3. As quark condensates determine the masses of baryons and quarks based on Eqs. (4) and (5), in Fig. 6 we then present their masses as order parameters for chiral phase transitions. Evidently, after quarkyonic phase transition takes place, the masses of $u$- and $d$-quarks decrease quickly with density and approach to their current masses, indicating the chiral phase transition with vanishing $\bar{n}_{u}^{s}$ and $\bar{n}_{d}^{s}$.
The baryon masses decrease with density as well, where the mass of $\Lambda$-hyperons becomes even smaller than $s$-quarks. This is mainly attributed to the structural function $\alpha_{S}$ ($<1$) adopted in this work, where three quarks form a bound state with large binding energies. The mass of $s$-quarks will decrease at much larger densities with vanishing $s$-quark condensate $\bar{n}_{s}^{s}$, which may be accompanied with a first-order deconfinement phase transition if we take $B=1$ GeV/fm3. The effects of Pauli blocking term in Eq. (4) become evident at large densities, where the masses of baryons increase drastically with $n_{\mathrm{b}}^{Q}$ if we adopt $B=1$ GeV/fm3. Consequently, as indicated in Fig. 5, baryons begin to dissolve with the corresponding number densities approaching zero.
Figure 7: Velocity of sound $v_{s}$ (upper panel), quark fraction $n_{\mathrm{b}}^{Q}/n_{\mathrm{b}}$ (middle panel) and pressure $P$ (lower panel) of compact star matter as functions of energy density, which are obtained adopting the parameters $f_{u}=f_{d}=f_{s}=0.5$ (black), 0.7 (red), 1 (blue) and $B=0$, 1 GeV/fm3.
In Fig. 7, we present the velocity of sound $v_{s}=\sqrt{\mbox{d}P/\mbox{d}E}$, quark fraction $n_{\mathrm{b}}^{Q}/n_{\mathrm{b}}$, and the corresponding EOSs predicted by the extended NJL model, where various combinations of parameters $f_{q}=0.5$, 0.7, 1 and $B=0$, 1 GeV/fm3 are adopted. At $P\lesssim 1$ MeV/fm3, a first-order core-crust phase transition takes place. Since the saturation properties of nuclear matter indicated in Table 2 resemble those of DD-LZ1 in RMF models, we adopt the DD-LZ1 EOS for neutron star crusts Xia et al. (2022a, b), where the core-crust transition pressure is fixed by Maxwell construction. This nevertheless should be improved in our future study, where the EOSs and structures for neutron star crusts should be fixed in the extended NJL model in a unified manner. At larger pressures, quarkyonic phase transition takes place with nonzero $n_{\mathrm{b}}^{Q}$, which is generally of third order except for the case with a first-order quarkyonic phase transition obtained by adopting $f_{q}=0.5$ and $B=0$. As density increases, quark condensates $\bar{n}_{u}^{s}$ and $\bar{n}_{d}^{s}$ eventually approach to zero at $P\gtrsim 10$ MeV/fm3, indicating the occurrence of chiral phase transitions. Most chiral phase transitions are of first order, which becomes a smooth crossover if we take $f_{q}=0.5$ and $B=0$ with a large quark fraction. As indicated in the upper and middle panels of Fig. 7, finally a deconfiement phase transition takes place with $n_{\mathrm{b}}^{Q}/n_{\mathrm{b}}=1$ if we adopt the parameter sets ($f_{q}$, $B$ in GeV/fm3): (0.5, 0), (0.5, 1), and (0.7, 0), where the EOSs are dominated by quark matter with $n_{u}\approx n_{d}\approx n_{s}$ and velocity of sound reduced. Meanwhile, within the density range indicated in Fig. 7, baryonic matter still dominates if we adopt the parameter sets (0.7, 1), (1, 0), and (1, 1), where the EOSs coincide with each other at large densities with a large velocity of sound.
Figure 8: Mass-radius relations of compact stars obtained with the EOSs presented in Fig. 7. The shaded regions indicate the mass-radius constraints from the binary neutron star merger event GW170817 LIGO Scientific and Virgo Collaborations (2018), PSR J0030+0451 and PSR J0740+6620 Riley et al. (2019, 2021); Miller et al. (2019, 2021).
Based on the EOSs presented in Fig. 7, we then examine the corresponding compact star structures by solving the Tolman-Oppenheimer-Volkov (TOV) equation
$\frac{\mbox{d}P}{\mbox{d}r}=\frac{(E+P)(M+4\pi r^{3}P)}{2Mr-r^{2}/G}\ \text{and }\frac{\mbox{d}M}{\mbox{d}r}=4\pi Er^{2}$ (25)
with the gravity constant $G=6.707\times 10^{-45}\ \mathrm{MeV}^{-2}$. The obtained $M$-$R$ relations of compact stars are then presented in Fig. 8. Various observational constraints on compact star structures are indicated by the shaded areas, i.e., the masses and radii fixed by pulse profile modeling of PSR J0030+0451 and PSR J0740+6620 Riley et al. (2019, 2021); Miller et al. (2019, 2021) as well as the gravitational wave measurements from of the binary neutron star merger event GW170817 LIGO Scientific and Virgo Collaborations (2018). Additionally, there are precise mass measurements for PSR J1614-2230 (1.928 $\pm$ 0.017 $M_{\odot}$onseca et al. (2016) and PSR J0348+0432 (2.01 $\pm$ 0.04 $M_{\odot}$ntoniadis et al. (2013) by analyzing the orbital motion of pulsars in a binary system Lattimer (2012), which is not indicated in Fig. 8 since PSR J0740+6620 also reaches the two-solar-mass constraint.
The $M$-$R$ relations of compact stars predicted by the extended NJL model adopting the parameter sets (0.7, 1), (1, 0), and (1, 1) agree well with the observational constraints, while the radii for two-solar-mass compact stars predicted by the parameter sets (0.5, 1) and (0.7, 0) lie in the lower ends of the PSR J0740+6620 constraints Riley et al. (2021); Miller et al. (2021). Additionally, if the parameter set (0.5, 0) is adopted, the radii of the corresponding compact stars become too small according to the observational constraints, suggesting that the first-order quarkyonic phase transition at small densities is in tension with pulsar observations.
IV Conclusion
In this work, we propose an extended NJL model to describe baryonic matter, quark matter, and their transitions in a unified manner. The baryons are treated as clusters made of three quarks, while a density-dependent structural function $\alpha_{S}$ is introduced to modulate the four(six)-point interaction strengths to reproduce the baryon masses in vacuum and medium Bentz and Thomas (2001); Reinhardt and Weigel (2012); Xia et al. (2014). Additional vector interactions are introduced using vector mesons $\omega$ and $\rho$, where the baryon-meson couplings are density-dependent and are fixed by reproducing nuclear matter properties Li and Han (2013); Oertel et al. (2017); Centelles et al. (2009); Brown (2013) and $\Lambda$-hyperon potential depth in nuclear medium Gal et al. (2016); Sun et al. (2016, 2018); Rong et al. (2021). A dampening factor $f_{q}\equiv f_{u}=f_{d}=f_{s}$ is introduced to modulates the coupling strengths between quarks and vector mesons, where in this work we take $f_{q}=0.5$, 0.7, and 1. Similar to the melting of light clusters in nuclear medium Typel et al. (2010), we have introduced a Pauli blocking term to baryon masses so that baryons eventually become unbound in the presence of a quark Fermi Sea Xia et al. ; Xia et al. (2023). Then the deconfinement phase transition can be viewed as a Mott transition of quark clusters Bastian et al. (2018).
Employing the extended NJL model, we then investigate the properties of dense stellar matter. As the density of nuclear matter increases, quarks emerge as quasi-free particles and coexist with baryons in the same volume. This phase is considered as quarkyonic matter, where quarks can not be exited to higher energy states in the presence of baryons due to confinement McLerran and Pisarski (2007); McLerran (2009); McLerran and Reddy (2019). Depending on the strengths of Pauli blocking term $B$ and quark-vector meson couplings $f_{q}$, both first-order and continues phase transitions are observed for quarkyonic, chiral, and deconfinement phase transitions. In particular, we find quarkyonic phase transition is generally of third order but becomes of first order if we adopt $f_{q}=0.5$ and $B=0$, which predicts a compact stars that are too compact according to observational constraints. Similar situation takes place for chiral phase transitions, which are generally of first order but becomes a smooth crossover if we take $f_{q}=0.5$ and $B=0$. At largest densities, a deconfinement phase transition takes place with baryons melted, where both first-order and continues transitions are identified. Adopting Maxwell construction for the first-order phase transitions, we then obtain the EOSs of dense stellar matter. The corresponding compact star structures are then fixed by solving the TOV equation, which are confronted with various astrophysical constraints LIGO Scientific and Virgo Collaborations (2018); Riley et al. (2019, 2021); Miller et al. (2019, 2021). Our prediction generally agrees well with observations, while the parameter set $f_{q}=0.5$ and $B=0$ predicts too small radii for compact stars.
Acknowledgements.
The author would like to thank Professors Bao-An Li, Toshiki Maruyama, Makoto Oka, Gubler Philipp, Ting-Ting Sun, Toshitaka Tatsumi and Nobutoshi Yasutake for fruitful discussions. This work was supported by the National Natural Science Foundation of China (Grant No. 12275234) and the National SKA Program of China (Grant No. 2020SKA0120300).